OpenStudy (anonymous):

Myra’s stamp collection consisted of 120 stamps in October. In March, her collection grew to 138 stamps. Find the percent increase between October and March. (Points : 1) 15% 9% 18% 12%

OpenStudy (anonymous):

Subtract the starting value from the current value. #1 138-120=18 difference Divide the result by the starting value. #2 120/18=.15 Multiply the result by 100 #3 .15*100=15%